Carla bought a pair of jeans for $48 when she went to pay for them, the sales clerk said she would receive $12 off. What percent is the discount? (Please solve in percent decrease formula, i got 25% but im not sure i was absent yesterday and this is my make up work)please explain because i dont get it PLEASE!!!!:)
Answer by stanbon(75887)

Carla bought a pair of jeans for $48 when she went to pay for them, the sales clerk said she would receive $12 off. What percent is the discount
---
Quantity discount: 12
Percent discount: 12/48 = 1/4 = 25%
======================================
Cheers,
Stan H.